The owner of a pizza parlor wants to make a profit of 80% of the cost for each pizza sold. If it costs $8.50 to make a pizza, at what price should each pizza be sold?

Answer by oberobic(2304)

80% of the cost is
.8 * 8.50 = 6.80
.
so the sales price should be the cost + required profit
8.50 + 6.80 = 15.30
.
or you could say he need to sell for 100% of cost + 80% of cost = 180% of the cost
1.8*8.50 = 15.30
.
Done.